The dimensions of a rectangle garden were 3 yards by 10 yard. When both dimensions are increased by equal amounts,the area of the garden doubled. Find the dimensions of the new garden
Answer by lwsshak3(11628)

The dimensions of a rectangle garden were 3 yards by 10 yard. When both dimensions are increased by equal amounts,the area of the garden doubled. Find the dimensions of the new garden
***
let x=amt both dimensions are increased by
area=length*width
original area=10*3=30
double the area=60
x+3=new dimension of width
x+10=new dimension of length
..
(x+3)(x+10)=60
x^2+13x+30=60
x^2+13x-30=0
(x+15)(x-2)=0
x=2
x+3=5
x+10=12
new dimension of width=5 yds
new dimension of length=12 yds